OpenStudy (anonymous):

Evaluate the expression when x = –12. 3x – 17 A. –53 B. –19 C. 19 D. 53

OpenStudy (anonymous):

just plug in - 12 for x

OpenStudy (anonymous):

3 x -12 - 17 = ?

OpenStudy (anonymous):

can you solve that?

OpenStudy (anonymous):

\[3\times(-12)-17\]

OpenStudy (anonymous):

\[-36 - 17 = ?\]

OpenStudy (anonymous):

-53 so A. is the answer :)

OpenStudy (anonymous):

You should close this question now and click best response.